# A Survey on Class-Agnostic Counting: Advancements from Reference-Based to Open-World Text-Guided Approaches

### v1

@@ Abstract
Object counting has recently shifted towards class-agnostic counting (CAC), which addresses the challenge of counting objects across arbitrary categories, tackling a critical need in versatile counting systems. While humans effortlessly identify and count objects from diverse categories without prior knowledge, most counting methods remain restricted to enumerating instances of known classes, requiring extensive labeled datasets for training, and struggling under open-vocabulary settings. Conversely, CAC aims to count objects belonging to classes never seen during training, typically operating in a few-shot setting. In this paper, for the first time, we review advancements in CAC methodologies, categorizing them into three paradigms based on how target object classes can be specified: reference-based, reference-less, and open-world text-guided. Reference-based approaches have set performance benchmarks using exemplar-guided mechanisms. Reference-less methods eliminate exemplar dependency by leveraging inherent image patterns. Finally, open-world text-guided methods utilize vision-language models, enabling object class descriptions through textual prompts, representing a flexible and appealing solution. We analyze state-of-the-art techniques and we report their results on existing gold standard benchmarks, comparing their performance and identifying and discussing their strengths and limitations. Persistent challenges – such as annotation dependency, scalability, and generalization – are discussed, alongside future directions. We believe this survey serves as a valuable resource for researchers to understand the progressive developments and contributions over time and the current state-of-the-art of CAC, suggesting insights for future directions and challenges to be addressed.
@@ II Existing Surveys, Scope, and Search Strategy
III Existing Approaches
III-A Problem Formulation
III-B Reference-based Class-agnostic Counting

### v2

@@ Abstract
Visual object counting has recently shifted towards class-agnostic counting (CAC), which addresses the challenge of counting objects across arbitrary categories – a crucial capability for flexible and generalizable counting systems. Unlike humans, who effortlessly identify and count objects from diverse categories without prior knowledge, most existing counting methods are restricted to enumerating instances of known classes, requiring extensive labeled datasets for training and struggling in open-vocabulary settings. In contrast, CAC aims to count objects belonging to classes never seen during training, operating in a few-shot setting. In this paper, we present the first comprehensive review of CAC methodologies. We propose a taxonomy to categorize CAC approaches into three paradigms based on how target object classes can be specified: reference-based, reference-less, and open-world text-guided. Reference-based approaches achieve state-of-the-art performance by relying on exemplar-guided mechanisms. Reference-less methods eliminate exemplar dependency by leveraging inherent image patterns. Finally, open-world text-guided methods use vision-language models, enabling object class descriptions via textual prompts, offering a flexible and promising solution. Based on this taxonomy, we provide an overview of the architectures of 29 CAC approaches and report their results on gold-standard benchmarks. We compare their performance and discuss their strengths and limitations. Specifically, we present results on the FSC-147 dataset, setting a leaderboard using gold-standard metrics, and on the CARPK dataset to assess generalization capabilities. Finally, we offer a critical discussion of persistent challenges, such as annotation dependency and generalization, alongside future directions. We believe this survey will serve as a valuable resource for researchers, highlighting the progressive developments in CAC and offering insights for future advancements in the field.
@@ II Existing Surveys, Scope, and Search Strategy
III Existing Approaches
III-A Problem Formulation
III-B Reference-based Class-agnostic Counting